​Arnold Tofu owns and operates a chain of vegetable protein "hamburger" restaurants in northern Louisiana. Sales figures and profits for the stores are in the table below.

Arnold Tofu owns and operates a chain of vegetable protein "hamburger" restaurants in northern Louisiana. Sales figures and profits for the stores are in the table below.

Store Profits Sales

1    14      6

2    11     3

3    15      5

4     16     5

5     24    15

6     28    18

Calculate a regression line for the data and determine the forecast of profit for a store with sales of $24.
Answer #1

Answer:


Given data,

Store

Sales(X)

Profits(Y)

XY

X2

Y2

1

6

14

84

36

196

2

3

11

33

9

121

3

5

15

75

25

225

4

5

16

80

25

256

5

15

24

360

225

576

6

18

28

504

324

784


Equation of least square regression line is: Y = a + b X

Now, Slope (b) = (NΣXY - (ΣX)(ΣY)) / (NΣX2 - (ΣX)2)

= (6 x 1136 - 52 x 108) / (6 x 644 - 522) = 1.034

So, b = 1.034

Now, Intercept(a) = (ΣY - b(ΣX)) / N

= (108-1.034x52) / 6 = 9.038

So, the regression line is Ft = 9.038 + 1.034 X

So, Sales is $24 million, profit = 9.038 + 1.034 X 24 = $ 33.85


answered by: Subrahmanyam golla